OpenStudy (anonymous):

i am having problems with integrals, the problem is (sec(x))^2 between pi/6 and -pi/6. i know that the integral tan(X) but when i use the first fundamental theorem of calculus i i get lost, i get the last step and i get lost help please

hartnn (hartnn):

so, you have \(\Large [\tan x]^{\pi/6}_{-\pi/6} = [\tan (\pi/6) - (-\tan (\pi/6))] = ...\)

OpenStudy (anonymous):

yes i have that, its the actual solving that i am having problems with

hartnn (hartnn):

you know whats tan (pi/6) = ... ?

hartnn (hartnn):

and i got that because \(\tan (-x) = - \tan x\)

OpenStudy (anonymous):

i do believe its .707?

OpenStudy (anonymous):

yeah i saw that

hartnn (hartnn):

tan (pi/6) = sin (pi/6) / cos (pi/6) = .... ?

OpenStudy (precal):

use the values from the unit circle

OpenStudy (anonymous):

its .6 which is rounded

hartnn (hartnn):

yes, or you can say its \(\Large \dfrac{1}{\sqrt 3}\)

hartnn (hartnn):

so, \(\Large [\tan x]^{\pi/6}_{-\pi/6} = [\tan (\pi/6) - (-\tan (\pi/6))] = ... \\ \Large 2 \tan (\pi/6) = \dfrac{2}{\sqrt 3}\)

OpenStudy (anonymous):

thank you, that helped, and very nice profile picture

hartnn (hartnn):

welcome ^_^ happy to help :) and since you're new here, \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)

OpenStudy (anonymous):

thanks lol